OpenStudy (anonymous):

Find equation of the plane through (-1,4,2) and containing the line of intersection of the planes 4x-y+z-2=0 , 2x+y-2z-3=0

OpenStudy (anonymous):

First do you know how to find the line of intersection of the 2 planes?

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

So what is it?

OpenStudy (anonymous):

we find the cross product of the normals of the two planes...and thats the vector parallel to normal of the two planes...then we find a point which lies on both the planes...???

OpenStudy (anonymous):

Hmm that's not how I learned it, you may be correct though. Actually yes that makes sense. This is how I would do it. Simply solve the matrix that is the solution set for both planes 4x-y+z=2 2x+y-2z=3 [4 -1 1 2] [2 1 -2 3] Row reduce that and you get [1 0 -1/6 5/6] [0 1 -5/3 4/3] The solution of this is: (5/6, 4/3, 0) + t*(1/6, 5/3, 1) =(5/6, 4/3, 0) + t*(1, 10, 6) You can try it your way and you should get the same thing

OpenStudy (anonymous):

OK now you have a vector on the plane (1,10,6) and you have 2 points: (-1,4,2) and (5/6, 4/3, 0). Any idea what to do next?

OpenStudy (anonymous):

okay i got till here...after that i just couldnt figure out what to do :/

OpenStudy (anonymous):

Why is your picture hitler??? Kind of offensive. Anyways, if you find 2 vectors that are on a plane, you can always find the equation for the plane by simply calculating their cross product. You have 1 vector, and you can get a second vector by simply finding the difference between the 2 points!

OpenStudy (anonymous):

Don't peak if you want to try it on your own! Here's my work: (-1,4,2) - (5/6, 4/3, 0) = (-11/6,8/3,2) (1,10,6) x (-11/6,8/3,2) = (4,-13,21) 4x-13y+21z = d 4(-1)-13(4)+21(2) = d -14 = d 4x-13y+21z = -14 :)

OpenStudy (anonymous):

Okay I have another question. If the plane was perpendicular to the line of intersection of the two planes... what would we do then? will the answer be the same as this one?

OpenStudy (anonymous):

How can a plane be perpendicular to a line?

OpenStudy (anonymous):

I don't know :/ I recently saw a question like that but can't find it now....

OpenStudy (anonymous):

A plane can't be perpendicular to a line, that makes no sense...

OpenStudy (anonymous):

oh i meant to say containing the line perpendicular to line of intersection of two planes

OpenStudy (anonymous):

sorry my bad

OpenStudy (anonymous):

That's fine! I just took linear last semester and I was SO confused. A plane can be perpendicular to a plane.

OpenStudy (anonymous):

You seem to be doing very well, this geometry stuff throws most people off but you're on track :)

OpenStudy (anonymous):

Oh you corrected yourself. I misread. "containing the line perpendicular to line of intersection of two planes". Hmm gimme a min

OpenStudy (anonymous):

Because there are infinite lines that are perpendicular to a line in 3 dimensions... I'm not sure sorry

OpenStudy (anonymous):

Okayy. Np. Thanks for your help. :)